OpenStudy (shamil98):

What is the Cartesian equation equivalent to the polar equation r = sec Ө csc Ө? x = r cos Ө y = r sin Ө x/r =cos Ө y/r = sin Ө r^2 = x^2 + y^2 and then i get lost. halp pls

zepdrix (zepdrix):

\[\Large\rm r=\sec \theta \csc \theta\]Do you remember how secant and cosecant relate to sine and cosine? :) That will help us quite a bit here.

OpenStudy (shamil98):

Yeah, sec = 1/cos csc = 1/sin

zepdrix (zepdrix):

\[\Large\rm r=\color{royalblue}{\sec \theta}\cdot\csc \theta\]Mmm ok that's a good start. \[\Large\rm r=\color{royalblue}{\frac{1}{\cos \theta}}\cdot\frac{1}{\sin \theta}\]

zepdrix (zepdrix):

If I divide both sides by r, I end up with:\[\Large\rm 1=\color{royalblue}{\frac{1}{r\cos \theta}}\cdot\frac{1}{\sin \theta}\]I'll divide by r again and get,\[\Large\rm \frac{1}{r}=\color{royalblue}{\frac{1}{r\cos \theta}}\cdot\frac{1}{r\sin \theta}\]

zepdrix (zepdrix):

Maybe that wasn't extremely necessary, I just like to try and work with those first two identities that you listed :)

zepdrix (zepdrix):

Do you see where we can plug them in?

OpenStudy (shamil98):

Not really, my brain is super-fried atm :(

zepdrix (zepdrix):

\[\Large\rm \frac{1}{r}=\frac{1}{\color{orangered}{r\cos \theta}}\cdot\frac{1}{\color{blue}{r\sin \theta}},\qquad \color{orangered}{x=r \cos \theta},\qquad \color{blue}{y=r \sin \theta}\]Ya? :d

OpenStudy (shamil98):

oh!

OpenStudy (shamil98):

so then r = r cos θ * r sin θ r = x*y

zepdrix (zepdrix):

r = xy ya looks good :) So then undo your r using the square identity thing, ya?

zepdrix (zepdrix):

\[\Large\rm r^2=x^2+y^2\]Therefore,\[\Large\rm r=\sqrt{x^2+y^2}\]

zepdrix (zepdrix):

We can plug that in :d

OpenStudy (shamil98):

so xy = sqrt(x^2 +y^2)

OpenStudy (shamil98):

and therefore y = sqrt(x^2+y^2)/x

zepdrix (zepdrix):

Well, we'd like to get it in the form \(\Large\rm y=x~stuff\). See how you left a y^2 on the right?

OpenStudy (shamil98):

oh.

zepdrix (zepdrix):

So let's back up a step,\[\Large\rm xy=\sqrt{x^2+y^2}\]and square each side.

OpenStudy (shamil98):

x^2 * y^2 = x^2 + y^2

zepdrix (zepdrix):

Ok good. Do you understand how to get your y's grouped up and isolated? It's a tricky couple of algebra steps I guess.

OpenStudy (shamil98):

Yeah, i'm gonna need some guidance on that part as well. x^2 * y^2 - y^2 = x^2 no clue what next lol

zepdrix (zepdrix):

Each term contains a y^2, let's factor a y^2 out of each :)\[\Large\rm y^2(\qquad -\qquad)=x^2\]What's going to be left in the brackets when we do that?

OpenStudy (shamil98):

ofc. y^2(x^2 -1) = x^2 so y^2 = (x^2-1)/x^2 y = sqrt( (x^2-1)/x^2))

OpenStudy (shamil98):

well +/- the sqrt of that

OpenStudy (shamil98):

Oh, i had it backwards its x^2 over that

zepdrix (zepdrix):

Ah yes good :) The original polar representation, if you graph it, it gives you 4 "branches". Since we now have a function, we need both the plus and minus root, each gives us 2 "branches".

zepdrix (zepdrix):

\[\Large\rm y=\pm\sqrt{\frac{x^2}{x^2-1}}\]Yay good job! \c:/

OpenStudy (shamil98):

thanks a lot!!! :)